what percentage of 300 is 40?

Answer by stanbon(75887)

what percentage of 300 is 40?
-----------------
Percentage means per 100
---
x/100 = 40/300
x = 100(40/300)
x = 40/3
x = 13 1/3
----------------------
40 is 13 1/3 % of 400
==========================
Cheers,
Stan H.